OpenStudy (anonymous):

Lim [(1/x-1)-(2/x^2-1)], x->1 Right answer is 1/2

OpenStudy (freckles):

combine fractions

OpenStudy (anonymous):

\[\frac{1}{x-1}-\frac{2}{x^2-1}=\frac{1}{x+1} \]